OpenStudy (tia_sweetheart17):

solve for y. x-9y=10

OpenStudy (mathstudent55):

To isolate y, you need to get rid of the x-part on from left side.

OpenStudy (mathstudent55):

Since there is x on the left side, you need to subtract x from both sides. Can you do that?

OpenStudy (anonymous):

x= 10 + 9y x-10 = 9y (x -10)/9 = y

OpenStudy (tia_sweetheart17):

does y=9 @mathstudent55

OpenStudy (mathstudent55):

x - 9y = 10 Subtract x from both sides: -9y = 10 - x Now divide both sides by -9: y = (10 - x)/(-9) Now simplify the right side: y = (x - 10)/9 \(y = \dfrac{x - 10}{9} \)

OpenStudy (mathstudent55):

That is the final answer.